OpenStudy (anonymous):

Create your own quadratic equation and demonstrate how it would be solved by graphing, factoring, the quadratic formula, and by completing the square

OpenStudy (anonymous):

you can use y=x^2+5x+6 now you can factor them

OpenStudy (anonymous):

y=(x+2)(x+3)=0 therefore x=-2,-3

OpenStudy (anonymous):

using the quadratic formula x=(-b-+sqrt(b^2-4ac))/2a

OpenStudy (anonymous):

x=(-b-+sqrt(b^2-4ac))/2a x=-5-+sqrt(5^2-4(1)(5)))/2(1) x=-2, -3

OpenStudy (anonymous):

graphing them if y=0, x=-2 and x=3 if x=0, y=6 also you can try plotting them using the software in the internet

OpenStudy (anonymous):

graphing them if y=0, x=-2 and x=3 if x=0, y=6 also you can try plotting them using the software in the internet

OpenStudy (anonymous):

hope that help.....

OpenStudy (anonymous):

@mark_0 how would you solve it by completing the square?